OpenStudy (anonymous):

An object that is translated 5 units to the right would be written as y + 5?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

like in an equation it has to be like y= x-5

OpenStudy (anonymous):

false, that would make the function move down 5

OpenStudy (anonymous):

how should the equatio be written then?

OpenStudy (anonymous):

do you have an equation for your function?

OpenStudy (anonymous):

y= a(x-5)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im just asking too since im not sure

OpenStudy (cwrw238):

if for example you moved the line y=x 5 units to the write the equation would then be y = x - 5

OpenStudy (anonymous):

moving a line down 5 is the same as moving it right 5, but quadratics, cubics, etc would be different

OpenStudy (anonymous):

^cool, I never actually saw that till now lol

OpenStudy (anonymous):

(x+5)

OpenStudy (cwrw238):

moving y = x^2 5 units to the right gives y = (x-5)^2